A pinball machine launch ramp is constructed by placing a spring of force constant ? at the bottom of a ramp of length ? inclined at 30 degree. Suppose the ramp is now waxed, so there is is no friction. What is the distance the spring should be compressed so that the ball just makes it to the top of the ramp?
Answer #1

Component of Force responsible for the motion, = F \cos \theta

Work done by this force, W= F \cos \theta * L = FL \cos \theta

This work is supplied by Spring, W = \frac{1}{2}kx^2

\Rightarrow \frac{1}{2}kx^2 =FL \cos \theta

The spring must be compressed by, \Rightarrow x =\sqrt{\frac{2FL \cos \theta}{k} } so that the ball just makes it to the top of the ramp